# MEMO: Dataset and Methods for Robust Multimodal Retinal Image Registration with Large or Small Vessel Density Differences

### v1

@@ 3.3 OCTA
The procedure for OCTA image acquisition is also shown in Fig. 3. OCTA scans centered on the fovea were taken using the same Heidelberg Spectralis with a $10\times 10$ degree protocol, consisting of 512 a-scans $\times$ 512 b-scans with 5-10 microns between b-scans and 5-7 frames averaged per b-scan location. Projection images of the superficial vascular plexus (SVP), intermediate capillary plexus (ICP), and deep capillary plexus (DCP) were generated using the segmentation algorithms and slab definitions provided by the Spectralis software (Heidelberg Eye Explorer, version 1.10.3.0, Heidelberg Engineering, Germany). The SVP slab was defined as between the internal limiting membrane to the anterior border of the inner plexiform layer, the ICP included the entire inner plexiform layer, and the DCP ranged from the posterior border of the inner plexiform layer to the anterior border of the outer plexiform layer. The projection images were processed using projection artifact removal (PAR).
All images from the OCTA image groups were stored in TIF format. Each of the OCTA image groups contained three images from the three layers (i.e., SVP, ICP and DCP). Fifteen OCTA image groups had the resolution of $512\times 512$ pixels, while the other 15 had a resolution of $768\times 768$ pixels. The SVP image from each OCTA image group was used for registration ground truth labeling.
@@ 4.1.1 LVD-Seg Background
As discussed in Section 2.2, vessel segmentation has been frequently used as the first step for multimodal retinal image registration [25, 27, 28, 47, 46, 48], primarily because features of vessels are considered to be more consistent across different modalities. Recently, deep learning-based vessel segmentation methods have shown superior performance. They can be categorized into two groups, supervised [25, 54] and unsupervised methods [27, 28], which present different limitations. The supervised vessel segmentation methods [25, 54] usually require a large number of high-quality pixel-level vessel masks for training to ensure test performance. However, such high-quality pixel-level vessel masks are often difficult and time-consuming to acquire. To avoid the need for pixel-level ground truth, an unsupervised vessel segmentation method based on style transfer has been proposed [27, 28]. However, because of the lack of direct supervision, we found that training the segmentation networks with style loss alone produced unreliable results on our MEMO dataset due to the large VD difference between EMA and OCTA images (Sec. 7.1). To alleviate the limitations of both supervised and unsupervised vessel segmentation methods, we designed a novel two-stage semi-supervised learning framework, LVD-Seg, to train our vessel segmentation module. Details of the two stages are described as follows.
@@ 4.1.2 LVD-Seg Stage 1 - Supervised Loss
In this stage, we trained our vessel segmentation module on $n$ manually-annotated EMA vessel segmentation masks, where $n$ could be as few as three according to our experiment results (Sec. 7.2). We used EMA vessel segmentation masks because the VD of EMA is much smaller than that of OCTA. This is because OCTA images contain a plethora of small capillaries which do not present in the corresponding EMA images and are not helpful for image registration. Moreover, labeling the less complex EMA vessel segmentation masks is much more feasible in terms of efficiency than labeling OCTA vessel segmentation masks.
@@
$I$ represents the input EMA image and $M$ represents the ground truth EMA mask. $Pred(I)$ represents the predicted segmentation mask of $I$. $i$ represents the $i^{th}$ pixel of the predicted segmentation mask or the ground truth mask. $N$ denotes the total number of pixels. In addition to MSE, the self-comparison loss [27, 28], denoted as $L_{sc}$, was also adopted to make the prediction robust against input image rotation. Specifically, $L_{sc}$ is defined as
$L_{sc}=MSE(Rot_{-90}(Pred(Rot_{90}(I)),Pred(I)),$ (2)
where $Rot_{\theta}(I)$ represents $I$ rotated by $\theta\degree$. Here, $L_{sc}$ can be seen as an alternative way to perform data augmentation. Overall, the training loss for stage 1, denoted as $L_{s1}$, can be written as
@@
where $w_{v}$ and $w_{sc}$ represent the weighting factors for the MSE and the self-comparison loss. In this paper, $w_{v}$ and $w_{sc}$ were set to 1 and 1e-3, respectively. The trained weights of the EMA vessel segmentation network in stage 1 were used to initialize both the EMA and OCTA vessel segmentation networks in stage 2.
@@ 4.1.3 LVD-Seg Stage 2 - Unsupervised Loss
Training a vessel segmentation network directly with supervised losses on very few ground truth vessel masks may not work well on a larger test set. Moreover, it is unknown whether the segmentation network trained on EMA images can extract vessels well in OCTA images for multimodal image registration, especially when a relatively large VD difference exists between the two modalities. To deal with these potential issues, we further optimized the segmentation networks using style loss [27, 28, 33] with a joint style target mask, a stand-alone EMA ground truth segmentation mask. Using a joint style target mask with style loss encouraged the EMA and OCTA segmentation networks to segment shared vessels.
Style loss penalized the difference between the predicted segmentation mask and the style target mask using Gram matrices. The Gram matrix was used to capture the style information but remove the spatial information, which can be written as
$G_{j}(I)=\frac{1}{C_{j}H_{j}W_{j}}\sum_{h=1}^{H_{j}}\sum_{w=1}^{W_{j}}\phi_{j}(I)\phi_{j}^{{}^{\prime}}(I).$ (4)
Here, $\phi_{j}(I)$ denotes the feature map with shape $C_{j}\times H_{j}\times W_{j}$ obtained from the $j^{th}$ layer of the pre-trained VGG-16 network [50] by feeding the network with input image $I$. Style loss ($L_{st}$) is then defined as the squared Frobenius norm of the difference between the Gram matrices of the predicted segmentation mask ($Pred(I)$) and the style target mask ($M_{t}$), which can be written as
@@
Note that the two different modalities used different weighting factors for style loss. The self-comparison loss was also adopted as an additional constraint for the predicted segmentation masks. In this paper, $w_{st}^{e}$, $w_{st}^{o}$ and $w_{sc}$ were set to 100, 1 and 1e-3, respectively.
@@ 4.2.1 Feature Detection and Description
We adopted pre-trained SuperPoint [55] as our feature detector and descriptor since it demonstrated good performance on detecting feature points of binary segmentation masks [27]. SuperPoint contains a shared encoder, an interest point decoder and a descriptor decoder. It was first trained on a synthetic dataset with labeled interest points to detect feature points. Next, Homographic Adaptation was used to self-label a large unlabeled real world image dataset. Finally, the model was jointly-trained to extract feature points and their corresponding descriptors with self-supervision. We refer the readers to [55] for more details. In this paper, the non-maximum suppression distance was set to 4 and the detector confidence threshold was set to 0.015 for keypoint detection.
@@ 4.2.2 Feature Matching and Registration
We determined the matched feature points based on bidirectional calculation of the minimum Euclidean distance. That is, a feature point X in an OCTA image is said to match a feature point Y in the corresponding EMA image only when the Euclidean distance between each other’s feature descriptors is smaller than (1) the Euclidean distance between X and any other feature point in the EMA image and (2) the Euclidean distance between Y and any other feature point in the OCTA image. The Random Sample Consensus (RANSAC) [56] method was applied to remove outliers and estimate the partial affine transformation matrix between the EMA and OCTA pair. Here, the partial affine transformation (i.e., 4 degrees of freedom) was adopted because the EMA and OCTA images in MEMO already have the same pixel density (i.e., scale factor).
@@ 5.1.1 The MEMO Dataset
The MEMO dataset contains 30 pairs of images. Fifteen pairs (with even indices) were selected as the training set, and the rest of the pairs (with odd indices) were used as the test set. For OCTA, the SVP layer projection images were used in our experiments because they contained clearer arterioles and venules, which could also be observed in EMA images. For EMA, the stacked image of each EMA image sequence was used for the purpose of denoising. Furthermore, we annotated the vessel segmentation mask of each EMA stacked image for the proposed VDD-Reg and MSD. We also annotated one EMA stacked image that is not part of the MEMO dataset as the style target. Note that even though we annotated the vessel masks for all EMA images, our VDD-Reg actually required only three of those to maintain its performance.
For the image pre-processing, the OCTA images were first resized to $256\times 256$ pixel. Then, the EMA images were resized using the same scaling factors of the corresponding OCTA images. Next, to meet the requirement of our model, the resized EMA images were then cropped to ensure that their widths and heights were multiples of 8. Finally, the annotated EMA vessel segmentation masks and the registration ground truth were pre-processed accordingly to ensure their correct scale. To ensure the quality and consistency of annotation, all annotations were drawn by the same human annotator and checked by an experienced ophthalmologist (OJS).
@@ 5.1.2 The CF-FA Dataset
The CF-FA dataset contains 59 pairs of color fundus (720 $\times$ 576, RGB) and fluorescein angiography (720 $\times$ 576, grayscale) images. We manually labeled 6 pairs of corresponding points for all image pairs as the registration ground truth. We selected 29 image pairs (with odd indices) as the training set, 29 image pairs (with even indices) as the test set, and 1 image pair (normal/1-1) as the style target. Similar to the MEMO dataset, we manually annotated the vessel segmentation masks of the style target and three selected color fundus images from the training set for the proposed VDD-Reg.
5.2 Baseline Methods
Table 2: Summary of the Baseline Methods
Methods Deep learning-based Indirect Transfer Method
SURF-PIIFD-RPM [43] - - -
LoFTR [57] - -
SG [58] - -
$*{}^{*}$ [29] CycleGAN
$*{}^{*}$ [28] Vessel Segmentation
VDD-Reg (Ours) Vessel Segmentation
Deep learning-based indicates whether the method includes any deep neural network. Indirect indicates whether the method applies a transfer method. $*{}^{*}$ indicates that the official implementation is not available, so we implement the method on our own. Some differences might exist, which are discussed in Section 5.2.
We compared VDD-Reg with five baseline methods listed in Table 2. For fair comparison, we made all methods except for SURF-PIIFD-RPM estimate the partial affine transformation matrix and adopt RANSAC with the same hyperparameters. For SURF-PIIFD-RPM, as we directly adopted the official code, the affine transformation matrix was applied and RANSAC was not used. Moreover, for methods that used SuperPoint [55] for keypoint detection and description, including SG, CycleGAN, Content-Adaptive and VDD-Reg, the official pretrained network was adopted without fine-tuning. More details about the baseline methods are listed as follows:
@@ 5.2 Baseline Methods
  • Content-Adaptive [28]: This method designed a content-adaptive vessel segmentation network based on the pixel-adaptive convolution (PAC) [51] guided by the phase images. The network was trained with style loss (Eq. 6) and the self-comparison loss (Eq. 2). The image registration loss based on the ground truth transformation matrix was also used. We implemented this method based on the code from [61]. Unlike the original paper [28], we ignored the outlier rejection network and did not fine-tuned SuperPoint because these were general techniques applicable to all the other competing methods, which were not within the scope of this paper.
@@ 5.3 Training and Testing Details
All networks in our method were implemented in PyTorch. For the vessel segmentation module, both stages took 1000 epochs for training. The trained networks in stage 1 were used to initialize the networks in stage 2. The Adam optimizer [62] with learning rate 1e-4 was used. A batch size of 1 was used due to the limitation of our GPU memory. OpenCV was used for data pre-processing and the RANSAC algorithm, where $cv2.estimateAffinePartial2D()$ was adopted by setting the maximum reprojection error to 5 pixels and the maximum iteration to 2000.

### v2

@@ 3.3 OCTA
OCTA scans centered on the fovea were taken using the same Heidelberg Spectralis with a $10\times 10$ degree protocol, consisting of 512 a-scans $\times$ 512 b-scans with 5-10 microns between b-scans and 5-7 frames averaged per b-scan location. Projection images of the superficial vascular plexus (SVP), intermediate capillary plexus (ICP), and deep capillary plexus (DCP) were generated using the segmentation algorithms and slab definitions provided by the Spectralis software (Heidelberg Eye Explorer, version 1.10.3.0, Heidelberg Engineering, Germany). The SVP slab was defined as between the internal limiting membrane to the anterior border of the inner plexiform layer, the ICP included the entire inner plexiform layer, and the DCP ranged from the posterior border of the inner plexiform layer to the anterior border of the outer plexiform layer. The projection images were processed using projection artifact removal (PAR).
All images from the OCTA image groups were stored in TIF format. Each of the OCTA image groups contained three images from the three layers (i.e., SVP, ICP and DCP). Fifteen OCTA image groups had the image size of $512\times 512$ pixels, while the other 15 had a image size of $768\times 768$ pixels. The SVP image from each OCTA image group was used for registration ground truth labeling.
3.4 Dataset Analysis
Despite its size, our MEMO dataset has sufficient diversity. To showcase this, additional image pairs corresponding to each eye of each non-human primate (NHP) are shown in Fig. 4. The images exhibit noticeable differences from each other, even if captured from the same NHP or eye. These differences arise from varying physical conditions of the NHPs, different capture times, imaging variations, etc. Besides, we present additional statistics of our MEMO dataset in Fig. 5, particularly focusing on the image registration aspect. It is observed that our MEMO dataset covers various image transformations.
Figure 4: Image samples corresponding to each eye of each NHP. The EMA image is placed on top of the OCTA image for each image pair.
Figure 5: The Statistics of our MEMO dataset. The number of image pairs (count) falling within different ranges of (a) translation in the x-axis (pixel), (b) translation in the y-axis (pixel), or (c) rotation (degree) are presented. The division of training and test data for the MEMO dataset is outlined in Sec. 5.1.1.
@@ 4.1.1 LVD-Seg Background
As discussed in Section 2.2, vessel segmentation has frequently been used as the first step for multimodal retinal image registration [25, 27, 28, 47, 46, 48], primarily because features of vessels are considered to be more consistent across different modalities. Recently, deep learning-based vessel segmentation methods have shown superior performance. They can be categorized into two groups, supervised [25, 55] and unsupervised methods [27, 28], which present different limitations. The supervised vessel segmentation methods [25, 55] usually require a large number of high-quality pixel-level vessel masks for training to ensure test performance, but such high-quality pixel-level vessel masks are often difficult and time-consuming to acquire. To avoid the need for pixel-level vessel masks, the unsupervised vessel segmentation methods based on style transfer have been proposed [27, 28]. However, due to the lack of direct supervision, the unsupervised vessel segmentation methods generally performs worse than the supervised ones in terms of the segmentation quality.
Unlike general vessel segmentation, which aims to accurately extract every vessel, the goal of the vessel segmentation module in our VDD-Reg is to extract vessels that are visible in both modalities. This is particularly crucial for multimodal retinal image registration when a majority of vessels in one modality (e.g. OCTA) are not visible in the other modality (e.g. EMA) due to the fundamental differences between the two modalities. To this end, we designed a novel two-stage semi-supervised learning framework, LVD-Seg, to train our vessel segmentation module. LVD-Seg was designed based on two key insights. First, LVD-Seg combined supervised and unsupervised training so that the resulted segmentation masks could be effectively used by the following registration module despite having very few pixel-level vessel masks. Second, only the pixel-level vessel masks from the modality with lower vessel density (e.g., EMA) were used as the supervisory signal for both supervised and unsupervised training, guiding the segmentation module to extract vessels that are visible in both modalities. Details of the two stages in LVD-Seg are described as follows.
@@ 4.1.2 LVD-Seg Stage 1 - Supervised Loss
In this stage, we trained our vessel segmentation module on $n$ manually-annotated EMA vessel segmentation masks, where $n$ could be as few as three according to our experiment results (Sec. 7.2). We used EMA vessel segmentation masks because the vessel density of EMA is much lower than that of OCTA and vessels that can be captured by EMA are visible in both modalities. In other words, OCTA images contain a plethora of small capillaries which do not present in the corresponding EMA images and are not helpful for image registration. Moreover, labeling the less complex EMA vessel segmentation masks is much more feasible in terms of efficiency than labeling OCTA vessel segmentation masks.
@@
where $I$ represents the input EMA image and $M$ represents the ground truth EMA mask. $Pred(I)$ represents the predicted segmentation mask of $I$. $i$ represents the $i^{th}$ pixel of the predicted segmentation mask or the ground truth mask. $N$ denotes the total number of pixels. In addition to MSE, the self-comparison loss [27, 28], denoted as $L_{sc}$, was also adopted to make the prediction robust against input image rotation. Specifically, $L_{sc}$ is defined as
$L_{sc}=MSE(Rot_{-90}(Pred(Rot_{90}(I))),\ Pred(I)),$ (2)
where $Rot_{\theta}(I)$ represents $I$ rotated by $\theta$ degree. Here, $L_{sc}$ can be seen as an alternative way to perform data augmentation. Overall, the training loss for stage 1, denoted as $L_{s1}$, can be written as
@@
where $w_{v}$ and $w_{sc}$ represent the weighting factors for the MSE and the self-comparison loss. In this paper, $w_{v}$ and $w_{sc}$ were set to 1 and 1e-3, respectively. Here, $w_{sc}$ was set to a relatively smaller value compared to $w_{v}$ to ensure that $L_{v}$, the primary supervisory signal in this stage, could effectively guide the learning process. The trained weights of the EMA vessel segmentation network in stage 1 were used to initialize both the EMA and OCTA vessel segmentation networks in stage 2.
@@ 4.1.3 LVD-Seg Stage 2 - Unsupervised Loss
To ensure the extraction of common vessels visible in both modalities, we further optimized the segmentation networks using style loss [27, 28, 33] with a joint style target mask, a stand-alone EMA ground truth segmentation mask. Using a joint style target mask with style loss guided the EMA and OCTA segmentation networks to segment shared vessels, improving the registration accuracy. Moreover, using the same trained weights to initialize both the EMA and OCTA vessel segmentation networks further stabilized the unsupervised training in this stage.
Style loss penalized the difference between the predicted segmentation mask and the style target mask using Gram matrices [57, 33]. The Gram matrix was used to capture the style information but remove the spatial information, whose elements can be written as
$G_{j}(I)_{c,c^{{}^{\prime}}}=\frac{1}{C_{j}H_{j}W_{j}}\sum_{h=1}^{H_{j}}\sum_{w=1}^{W_{j}}\phi_{j}(I)_{h,w,c}\phi_{j}(I)_{h,w,c^{{}^{\prime}}}.$ (4)
Here, $\phi_{j}(I)$ denotes the feature map with shape $C_{j}\times H_{j}\times W_{j}$ obtained from the $j^{th}$ layer of the pre-trained VGG-16 network [50] by feeding the network with input image $I$, and $c,c^{{}^{\prime}}\in[1,\ C_{j}]$. The Gram matrix, a $C_{j}\times C_{j}$ matrix, is utilized to capture correlations among features, thus representing the ‘style’ of the input image $I$. Style loss ($L_{st}$) is then defined as the squared Frobenius norm of the difference between the Gram matrices of the predicted segmentation mask ($Pred(I)$) and the style target mask ($M_{t}$), which can be written as
@@
Note that the two different modalities used different weighting factors for style loss. The self-comparison loss was also adopted as an additional constraint for the predicted segmentation masks. In this paper, $w_{st}^{e}$, $w_{st}^{o}$ and $w_{sc}$ were set to 100, 1 and 1e-3, respectively. Similar to Eq. 3, $w_{sc}$ was set to a relatively smaller value to ensure that the primary supervisory signals in this stage, $L_{st}^{e}$ and $L_{st}^{o}$, effectively guided the learning process. Additionally, we observed that $L_{st}^{o}$, representing the style loss for OCTA, tended to be larger than $L_{st}^{e}$ due to the use of the EMA vessel segmentation ground truth mask as the style target. To balance the impacts of $L_{st}^{o}$ and $L_{st}^{e}$, we assigned $w_{st}^{e}$ to a relatively larger value compared to $w_{st}^{o}$.
@@ 4.2.1 Feature Detection and Description
We adopted pre-trained SuperPoint [56] as our feature detector and descriptor. It demonstrated superior performance than many traditional feature detectors and descriptors, and has been widely adopted in many applications which require feature matching, including multimodal image registration [27, 58, 59], localization [60] and two-view geometry estimation [61]. The network structure of SuperPoint is shown in Fig. 7. It contains a shared encoder, an interest point decoder and a descriptor decoder. It was first trained on a synthetic dataset with labeled interest points to detect feature points. Next, Homographic Adaptation [56] was used to self-label a large unlabeled real world image dataset. Finally, the model was jointly-trained to extract feature points and their corresponding descriptors with self-supervision. We refer the readers to [56] for more details. In this paper, the non-maximum suppression distance was set to 4 and the detector confidence threshold was set to 0.015 for keypoint detection.
Figure 7: The network structure of SuperPoint [56]. The green circles indicate the feature points detected by SuperPoint.
@@ 4.2.2 Feature Matching and Registration
We determined the matched feature points based on bidirectional calculation of the minimum Euclidean distance. That is, a feature point X in an OCTA image is said to match a feature point Y in the corresponding EMA image only when the Euclidean distance between each other’s feature descriptors is smaller than (1) the Euclidean distance between X and any other feature point in the EMA image and (2) the Euclidean distance between Y and any other feature point in the OCTA image. The Random Sample Consensus (RANSAC) [62] method was applied to remove outliers and estimate the partial affine transformation matrix between the EMA and OCTA pair. Here, the partial affine transformation (i.e., 4 degrees of freedom) was adopted because the EMA and OCTA images in MEMO were captured using the same device and already have the same pixel density (i.e., scale factor).
@@ 5.1.1 The MEMO Dataset
The MEMO dataset contains 30 pairs of images. Fifteen pairs (with even indices) were selected as the training set, and the rest of the pairs (with odd indices) were used as the test set. For OCTA, the SVP layer projection images were used in our experiments because they contained clearer arterioles and venules, and were free of projection artifact. For EMA, the stacked image of each EMA image sequence was used for the purpose of denoising. Furthermore, we annotated the vessel segmentation mask of each EMA stacked image for VDD-Reg. We also annotated one EMA stacked image that is not part of the MEMO dataset as the style target. Note that even though we annotated the vessel masks for all EMA images, our VDD-Reg actually required only three of those to maintain its performance.
For the image pre-processing, the OCTA images were first resized to $256\times 256$ pixels. Then, the EMA images were resized using the same scaling factors of the corresponding OCTA images. Next, to meet the requirement of our model, the resized EMA images were then cropped to ensure that their widths and heights were multiples of 8. Finally, the annotated EMA vessel segmentation masks and the registration ground truth were pre-processed accordingly to ensure their correct scale. To ensure the quality and consistency of annotation, all annotations were drawn by the same human annotator and checked by an experienced ophthalmologist (OJS).
@@ 5.1.2 The CF-FA Dataset
The CF-FA dataset contains 59 pairs of color fundus (720 $\times$ 576, RGB) and fluorescein angiography (720 $\times$ 576, grayscale) images. Twenty-nine pairs of images are from healthy subjects, while the other 30 pairs are from patients with retinopathy. We manually labeled 6 pairs of corresponding points for all image pairs as the registration ground truth. We selected 29 image pairs (with odd indices) as the training set, 29 image pairs (with even indices) as the test set, and 1 image pair (normal/1-1) as the style target. Similar to the MEMO dataset, we manually annotated the vessel segmentation masks of the style target and three selected color fundus images from the training set for the proposed VDD-Reg. For image pre-processing, the color fundus (CF) images were converted to grayscale, with no additional pre-processing steps applied.
5.2 Comparison to Existing Methods
Table 2: Summary of Five Existing Methods
Methods Year Deep learning-based Indirect Transfer Method
SURF-PIIFD-RPM [43] 2015 - - -
LoFTR [63] 2021 - -
SG [64] 2020 - -
CycleGAN-based [29] 2022 CycleGAN [52]
Content-Adaptive [28] 2021 Vessel segmentation
VDD-Reg (Ours) - Vessel segmentation
Deep learning-based indicates whether the method includes any deep neural network. Indirect indicates whether the method applies a transfer method. indicates that the official implementation is not available, so we implement the method on our own. Some differences might exist, which are discussed in Section 5.2.
We compared VDD-Reg with five existing methods listed in Table 2. These existing and previously described methods were selected to represent each type of methods we discussed in Sec. 2.2. SURF-PIIFD-RPM [43] is one of the well-known traditional methods for multimodal retinal image registration that demonstrated good performance. For the deep-learning based methods, two direct and two indirect methods were chosen for comparison. SuperGlue (SG) [64] and LoFTR [63] are two direct methods for feature detection and description which were proposed more recently. For indirect methods, we selected two methods [29, 28] that utilize different transfer methods, including CycleGAN [52] and vessel segmentation.
For fair comparison, we made all methods except for SURF-PIIFD-RPM estimate the partial affine transformation matrix and adopt RANSAC with the same hyperparameters. For SURF-PIIFD-RPM, as we directly adopted the official code, the affine transformation matrix was applied and RANSAC was not used. Moreover, for methods that used SuperPoint [56] for keypoint detection and description, including SG, CycleGAN-based, Content-Adaptive and VDD-Reg, the official pretrained network was adopted without fine-tuning. More details about these five existing methods are listed as follows:
@@ 5.2 Comparison to Existing Methods
  • Content-Adaptive [28]: This method designed a content-adaptive vessel segmentation network based on the pixel-adaptive convolution (PAC) [51] guided by the phase images. The network was trained with style loss and the self-comparison loss. The image registration loss based on the ground truth transformation matrix was also used. We implemented this method based on the code from [67]. Unlike the original paper [28], we ignored the outlier rejection network and did not fine-tuned SuperPoint because these were general techniques applicable to all the other competing methods, which were not within the scope of this paper.
@@ 5.3 Training and Testing Details
All networks in our method were implemented in PyTorch. The network architectures were implemented based on the implementations provided in [27] and [56], which were adopted with default settings for the DRIU and SuperPoint models, respectively. For the vessel segmentation module, both stages took 1000 epochs for training. The trained networks in stage 1 were used to initialize the networks in stage 2. The Adam optimizer [68] with learning rate 1e-4 was used. A batch size of 1 was used due to the limitation of our GPU memory. OpenCV was used for data pre-processing and the RANSAC algorithm, where the $cv2.estimateAffinePartial2D$ function, an OpenCV function that estimates partial affine transformation, was adopted by setting the maximum reprojection error to 5 pixels and the maximum iteration to 2000.